OpenStudy (anonymous):

2.12 + 42 + 72 + ... + (3n - 2)2 = quantity n times quantity six n squared minus three n minus one all divided by two

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

I have to state whether the statement is true or false for all positive integers n but Idk how to do this problem, it doesn't make sense

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

This?\[ 12+42+72+\ldots +(3n-2)^2=n\left(\frac{6n^2-3n-1}{2}\right) \]

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

But \((3(1)-2)^2 = (3-2)^2 = 1^2=1\)

OpenStudy (anonymous):

And \[ (3(2)-2)^2 = (6-2)^2 = 4^2=16 \]

OpenStudy (anonymous):

\[ (3(3)-2)^2 = (9-2)^2 = 7^2 = 49 \]

OpenStudy (anonymous):

good point none of these numbers are squares is there a typo somewhere?

OpenStudy (anonymous):

OH SORRY its supposed to be 1^2, 4^2, 7^2, ...

OpenStudy (anonymous):

Okay, so the first step is to let \(n=1\) and then simplify: \[ n\left(\frac{6n^2-3n-1}{2}\right) \]

OpenStudy (anonymous):

\[ 1\left(\frac{6(1)^2-3(1)-1}{2}\right) = \frac{6-3-1}{2} = \frac 22 = 1 \]So it works for \(n=1\).

OpenStudy (anonymous):

ahh ok I wrote my numbers wrong when I wrote the problem. so the statement would be true, correct?

OpenStudy (anonymous):

hmm is this supposed to be a proof by induction, or do you just need to check that \[\sum_{k=1}^n(3k-2)^2=n\left(\frac{6n^2-3n-1}{2}\right)\]

OpenStudy (anonymous):

\[ 2\left(\frac{6(2)^2-3(2)-1}{2}\right) = 6(4)-3(2)-1 = 24-6-1 = 17 \]So it doesn't seem to work for \(n=2\). Unless I made a mistake.

OpenStudy (anonymous):

Wait a moment

OpenStudy (anonymous):

\[1 +16 = 17\] So actually it does work for \(n=2\)

OpenStudy (anonymous):

so would the statement be true?